OpenStudy (anonymous):

Which of the following is equivalent to log (base 72) 9 ? log9 / log72 log72 / log9 log 8 log 1/8

OpenStudy (anonymous):

\[\bf \log_{a}(x)=\frac{ \log(x) }{ \log(a) }\]

OpenStudy (anonymous):

@EmilyFaye Which one do you think it is after looking at the above rule?

OpenStudy (anonymous):

b?